OpenStudy (anonymous):

Determine the value of “n” so that the given condition is satisfied: One root of \(16x^2 - 9x + n\) is three times the other.

OpenStudy (lgbasallote):

let x-k be one factor...the other would be x - 3k using remainder theorem p(k) = 0 16(k)^2-9(k) + n =0 16k^2 - 9k + n = 0 p(3k) = 0 16(3k)^2 - 9(3k) + n =0 16(9k^2) - 27k + n = 0 144k^2- 27k + n = 0 that's all i got lol

OpenStudy (anonymous):

lol...let's see what I can do now...

OpenStudy (anonymous):

144k^2 - 27k = 16x^2 - 9k 128k^2 - 18k = 0 k = 0 128k - 18 = 0 k = 9/64 Well, I just looked up the solutions...I was so close...

OpenStudy (anonymous):

I got 243/4096 and it turned out to be 243/256...interesting answer...

OpenStudy (anonymous):

Thanks lgba

OpenStudy (anonymous):

Oops posted wrong solution...

OpenStudy (lgbasallote):

lol i couldnt get *that* far...nice

OpenStudy (anonymous):